# A secure and private ensemble matcher using multi-vault obfuscated templates

### v1

@@ Abstract
Given the irrevocability of biometric samples and mounting privacy concerns, biometric template security and secure matching are among the essential features of any well-designed modern biometric system. In this paper, we propose an obfuscation method that hides the biometric template information with just enough chaff. The main idea is to reduce the number of chaff points to a practical level by creating $n$ sub-templates from the original template and hiding each sub-template with $m$ chaff points. During verification, $s$ closest vectors to the biometric query are retrieved from each vault and then combined to generate hash values that are compared with the stored hash value. We demonstrate the effectiveness of synthetic facial images, generated by a Generative Adversarial Network (GAN), as “random chaff points” within a secure-vault authorization system. This approach safeguards user identities during training and deployment. We tested our protocol using the AT&T, GT, and LFW face datasets, with the ROC areas under the curve being 0.99, 0.99, and 0.90, respectively. These numbers were close to those of the unprotected templates, showing that our method does not adversely affect accuracy.
@@ I Introduction
Biometric user identification has become a popular and convenient way for users to assert their identity across different devices and services, from cellular phones and laptops to access control and payment systems and even government services. However, many biometric systems are still vulnerable to certain attack vectors, such as sensor hijacks, injection of forged biometric templates, and network attacks on servers [1]. Such vulnerabilities may expose a biometric system to correlation attacks [2], replay attacks, Man-in-the-Middle attacks, Denial-of-Service attacks, or Brute-Force attacks [3, 4, 5]. Security incidents and privacy breaches through social media platforms, healthcare data, and government services [6, 7, 8] exemplify such vulnerabilities. Per the recent report from the HIPAA journal [9], there has been a substantial increase in data breaches across US hospitals, including data theft, unauthorized access, and hacking, highlighting the need for securing biomedical and biometric reference data.The same report also shows a large increase in the monetary penalties issued to hospitals due to such data breaches.
@@
To protect stored biometric references from such attacks, researchers have offered various security schemes such as random projection, feature disentanglement, fuzzy vaults (with helper data), and deep learning-based concealable multi-biometrics to mitigate the earlier-mentioned privacy and security issues [10, 11, 12, 13]. However, many of these approaches are either computationally too expensive or not completely secure against certain attack vectors. In this work, we propose an encrypted vault approach to securely store biometric templates. The reference feature vector, derived from the enrollment image and used for comparison with another incoming images, is called an enrollment template. For example, in the context of a Convolutional Neural Network (CNN) trained for face matching, this could be the flattened input to an FC layer. If the target image corresponds to the same person or object, for instance by the corresponding templates having a cosine similarity below a set threshold, the claimant is accepted. In this case the claimant is genuine, and the acceptance is a true positive of genuine accept. If the claimant is an impostor (the incoming image is not from the same identity) and they mange to pass the template comparison, the result is a false or impostor accept. A genuine comparison is anticipated to yield a higher cosine similarity, indicating that the feature vectors derived from the images closely match (Fig. 1). Our approach uses mutually exclusive embeddings generated by various deep learning models, combined and then securely encrypted using a SHA-512 based encryption. This technique also stores a large number of face embeddings generated from random identities (2000-4000 or higher), all stored in a vault, making it practically impossible to identify the original template. This attribute makes our approach robust against brute force attacks. We note that many other encryption-based approaches have been proposed previously to overcome such attacks [14, 15], but these approaches often adversely increase the time complexity of the biometric pipeline. Our approach does not inflict such a high computational burden, as we demonstrate later for the face verification use case, while it maintains the security of the process.
@@
The rest of this paper is organized as follows: section II focuses on related work. Section III describes our methodology and the main idea behind the framework, and explains the improvement in True Positive Rate (TPR) and the enhanced security of our approach. Section IV lays out the experimental setup. Section V details our results and their implications. Finally, we conclude by summarizing the unique attributes of our method, the difficulties faced, and future work.
@@ II Related Work
Many methods have been introduced over the years to preserve the privacy of biometric data. For instance, [16] used a de-identification technique. However, there was no guarantee of anonymity. Other methods tried encryption-based schemes, such as [17]. However, such solutions are often vulnerable to attack vectors such as the False Acceptance Rate (FAR) attack [18], brute force, and linkage attacks. Our approach can be shown to be secure against such attacks, as we demonstrate in Section III.B. Some authors used CNNs for protecting face templates while attaining good matching performance [19]. In another group of papers such as [20], researchers discuss blockchain-based methods for storing biometric templates. However, often execution time is higher for such schemes. Some studies have explored methods involving adversarial learning to conceal sensitive information within a learned space [21]. Specifically, there is a focus on disentangling identity from attribute information to mitigate soft-biometric information in face templates [22, 11]. Recent studies have investigated combining two biometric modalities, like fingerprint and iris scan, using various fusion techniques [23, 24]. While enhancing security through redundancy, this approach introduces a single point of failure. If one biometric trait is compromised or unavailable, the entire authentication process could be vulnerable.
Figure 1: Cosine similarity for template comparison
II-A Fuzzy Extractor
Fig. 2 describes a fuzzy extractor scheme and its terminology. We will use the same structure and terminology for our proposed obfuscation method. The Gen function generates a secret key and a public helper $P$ from a biometric template $t$. The Rep function takes a biometric query $q$ and the helper as its input. If dist($q,t$) is smaller than a preset threshold, then it generates the correct key. In other words, the helper $P$ helps to remove noise from query $q$ in order to generate $t$. The helper $P$ is required to be secure and reveal limited information about template $t$ to preserve the privacy of the template $t$. To satisfy the information-theoretic security, the entropy loss must be at most $\mathcal{L}$, as further detailed below [25],[26],[27].
@@ III Methodology
In this section, we will first informally explain the main idea behind the proposed method, and then provide the formalized algorithm. The idea is to divide the biometric template into $m$ sub-templates and hide them with $n$ chaff points to make it computationally impractical for an adversary to find the template. One solution would be to hide a template $t$ with $n$ chaff points in a public vault, $V=[r_{1},...,t,...,r_{n}]$ and store $f(t)$, where $f$ is a one-way function such as SHA-256. Now the question is how hard it would be for an adversary to find the index of $t$ in $V$. To make it computationally impractical for the adversary to find $t$, $n$ must be a very big number $(\lambda$ (security parameter)), as the best strategy to find the template $t$ is to compute $f(x)$ for all $x$ in the vault to find the template $t$. However, this method is also computationally impractical for the legitimate user. By way of example, assume we will have $f(t)$ and one vault containing $1+2^{80}$ vectors as public information (helper). We assume the inserted chaff points are indistinguishable from the template $t$ in the collection. By indistinguishable chaff points, we mean that, for example, if we were working with face templates stored in the vault, the chaff points would be representing (albeit different) faces in a way that an attacker wouldn’t be able to distinguish them from the actual template. For matching a query $q$, we compute the distance of $q$ from all points inside the vault and retrieve $m$ (depending on the threshold of the system) closest points to $q$: ${p_{1},p_{2},...,p_{m}}$, and compute $f(p_{1}),f(p_{2}),...,f(p_{m})$ and check if there are any matches between these values and $f(t)$. This system is secure under certain conditions, but impractical because during the authentication the user needs to compute the desired distance of the query $q$ vs. $1+2^{80}$ vectors.
@@
Thus we propose generating $n$ sub-templates $t_{1},t_{2},...,t_{n}$ from template $t$, using protocol $\mathcal{G}$, such that $\mathcal{G}(t):t_{1},t_{2},...,t_{n}$ and hide each of them with $m$ chaff points. In the end, we store $f([t_{1},t_{2},...,t_{n}])$ (, where $f$ is a one-way function), in $n$ vaults such that each vault has $m+1$ vectors. To make it computationally impractical for the adversary to find the template by brute-force, we need $m^{n}$ to be very large. Unlike the earlier impractical example, we can choose $m$ and $n$ to garner computational security against brute force attacks while having a practical solution. In the forthcoming experimental demonstration we use $m=2000$, and $n=5$, so that $m^{n}=2^{54.82}$ to get acceptable security against a brute force attack. When matching query $q$, we generate $q_{1},q_{2},...q_{n}$ using protocol $\mathcal{G}(q):q_{1},q_{2},...,q_{n}$ (protocol $\mathcal{G}$ will be explained later). The distances between $q_{i}$ and all $m+1$ vectors in the corresponding vault are then computed and the top $s$ vectors from each vault are retrieved. The threshold of the system is $s$ and the distance can be computed by any distance metric, such as cosine similarity. At this point, we have retrieved $sn$ vectors. Next we construct all $s^{n}$ possible combinations of vectors $v_{1},v_{2},...v_{s^{n}}$ and compute $f(v_{1}),f(v_{2}),...f(v_{s^{n}})$ and check if there are any matches between the stored value of $f([t_{1},t_{2},...,t_{n}])$ and the $f(v_{i})$. If $q$ is a genuine vector, it should be able to retrieve $t_{1},t_{2},...,t_{n}$ with the threshold $s$. If $q$ is an imposter, it will fail to retrieve $t_{1},t_{2},...t_{n}$ from the vaults. It should be noted that $s$ cannot be a large number as it would render the biometric matching computationally impractical.
III-A True Positive Rate (TPR) Improvement
@@ III-B The Algorithm
Sub-template Generator $\mathcal{G}(t,n)$: This protocol takes template $t$ and the number of vaults $n$ as its inputs and generates $n$ independent sub templates $T=(t_{1},t_{2},...,t_{n})$. In the experimental section, we propose one solution for such a function.
@@ IV Security
Revocability: A biometric system needs to be revocable because if it is compromised, or if the user wishes to no longer be identified, they should be able to revoke their reference. Our scheme’s templates are revocable by choosing a new set of chaff points and a new sub-template generator.
@@
Diversity: A secure template must not allow cross-matching across databases, thereby ensuring the user’s privacy. In our system, for a false positive to occur, the imposter must meet the authentication criteria of each vault and successfully retrieve all segments of the template, hence our system improves the cross-matching as compared to systems using one model. It also could be verified throughout the experimental section, the FPR is very small, because an imposter needs to satisfy multiple vaults to be able to positively match.
@@
Accuracy: Our method shows minimum accuracy degradation with a 0.99 area under the curve (AUC) for both AT&T database of faces [29] and Georgia Tech (GT) face database [30].
@@
Security: To show that our system is secure, we need to show that the helper data $P=(P_{1},P_{2})$ is not revealing too much information and there is enough entropy in this system. Chaff points are generated from the same distribution that the corresponding sub-template is generated. Hence we can say the sub-templates inside each vault are indistinguishable from other chaff points there. Getting this indistinguishability between the sub-template and chaff points inside a vault is straightforward; the challenge is to achieve indistinguishability between sub-templates and chaff points. In other words, if an adversary knows that $t_{1}$ is a sub-template, other $t_{i}$ should remain indistinguishable from chaff points, we achieve this by using independent models, given two models $M_{1},M_{2}$ and template $t$ and a random face $r$, let ======= = = be coming from the same image.
$|Pr[M_{1}(t)===M_{2}(t)]-Pr[M_{1}(t)===M_{2}(r)|=\epsilon$.
@@
To achieve this, we must make sure that the sub-templates are independent. If we assume everything looks totally random inside vaults, then we can claim the best strategy to find the template for any adversary is to check all possible combinations of elements in vaults, which we have made impractical by choosing proper $n$,$m$. As a result, our approach provides superior security against brute force attacks, as well as faster face acceptance and rejection capabilities.
@@
It is clear that given a face template $t$, a random face $r$, and two biometric models $M1$ and $M2$, $y_{1}=M1(t)$ and $y_{2}~{}=~{}M2(t)$ cannot be completely independent. However, in our case we are using big neural networks. With this assumption and if $M1$ and $M2$ have completely different structures, it is not going to be easy to distinguish between $M1(t)$ and $M1(r)$ given $M2(t)$. In other words, if we tell the adversary $M2(t)$ belongs to person $A$, and then if we ask him to tell us which one of $M1(t)$ and $M1(r)$ belongs to the same person, it will not be easy for the adversary to solve this problem.
@@
One approach is to reverse engineer the template to recreate the template and see which one of reverse ($M1(t))$ and reverse ($M1(r))$ are closer to the reverse ($M2(t)$). We have addressed this attack by adding randomness now the adversary needs to distinguish between $R_{2}M1(t)+R_{2}^{\prime}$ , $R_{2}M1(r)+R_{2}^{\prime\prime}$ , given $R_{1}M2(t)+R_{1}^{\prime}$.
V Experimental setup

### v2

@@ Abstract
Generative AI has revolutionized modern machine learning by providing unprecedented realism, diversity, and efficiency in data generation. This technology holds immense potential for biometrics, including for securing sensitive and personally identifiable information. Given the irrevocability of biometric samples and mounting privacy concerns, biometric template security and secure matching are among the most sought-after features of modern biometric systems. This paper proposes a novel obfuscation method using Generative AI to enhance biometric template security. Our approach utilizes synthetic facial images generated by a Generative Adversarial Network (GAN) as “random chaff points” within a secure vault system. Our method creates $n$ sub-templates from the original template, each obfuscated with $m$ GAN chaff points. During verification, $s$ closest vectors to the biometric query are retrieved from each vault and combined to generate hash values, which are then compared with the stored hash value. Thus, our method safeguards user identities during the training and deployment phases by employing the GAN-generated synthetic images. Our protocol was tested using the AT&T, GT, and LFW face datasets, achieving ROC areas under the curve of 0.99, 0.99, and 0.90, respectively. Our results demonstrate that the proposed method can maintain high accuracy and reasonable computational complexity comparable to those unprotected template methods while significantly enhancing security and privacy, underscoring the potential of Generative AI in developing proactive defensive strategies for biometric systems.
@@ I Introduction
Biometric user verification has become ubiquitous, but its vulnerabilities to various attacks and privacy breaches emphasize the need for enhanced security measures. Recent advances in Generative AI (GenAI) have transformed the field of machine learning, offering new opportunities for improving biometric security and privacy. GenAI’s ability to generate highly realistic and diverse synthetic data has been leveraged in various applications, including biometric security [17, 25]. When it comes to biometric security and privacy, GenAI may be used to synthesize biometric data for template obfuscation thus safeguarding user privacy. Unlike traditional methods, GenAI-produced chaff points can be highly varied and realistic, making it exceedingly difficult for attackers to distinguish between genuine and synthetic data, significantly enhancing biometric template security [49, 46, 8].
@@
Biometric systems have been vulnerable to various attack vectors, including sensor hijacks, injection of forged biometric templates, and network attacks on servers [34]. Such vulnerabilities may expose biometric systems to correlation attacks [53], presentation attacks, replay attacks, Man-in-the-Middle attacks, Denial-of-Service attacks, or Brute-Force attacks [20, 10, 28, 40]. Security incidents and privacy breaches through social media platforms, healthcare data, and government services [44, 36, 38] exemplify these vulnerabilities. The need for securing biomedical and biometric reference data is highlighted by the substantial increase in data breaches across US institutions and hospitals, including data theft, unauthorized access, and hacking, as reported by the HIPAA journal [22].
@@
To protect stored biometric references from such attacks, researchers have proposed various security schemes, such as random projection, feature disentanglement, fuzzy vaults with auxiliary (helper) data, and deep learning-based concealable multi-biometrics to mitigate the earlier-mentioned privacy and security issues [47, 3, 43, 1]. However, many of these approaches are either computationally too expensive or not completely secure against certain attack vectors. In this work, we propose an encrypted vault approach to securely store biometric templates. Our approach uses mutually exclusive embeddings generated by various deep learning models, combined and then securely encrypted using SHA-512-based encryption. This technique incorporates 2000-4000 AI-generated face embeddings as chaff points, stored alongside the genuine template in a secure vault. This makes it computationally infeasible to distinguish the original template from synthetic data. This attribute makes our approach robust against brute-force attacks.
The reference feature vector, derived from the enrollment image and used for comparison with other incoming images, is called an enrollment template. For example, in the context of a Convolutional Neural Network (CNN) trained for face matching, this could be the flattened input to an FC layer. If the target image corresponds to the same person or object, for instance by the corresponding templates having a cosine similarity below a set threshold, the claimant is accepted. In this case, the claimant is genuine, and the acceptance is a true positive or genuine accept. If the claimant is an impostor (the incoming image is not from the same identity) and they manage to pass the template comparison, the result is a false or impostor accept. A genuine comparison is anticipated to yield a higher similarity, indicating that the feature vectors derived from the images closely match.
Fig. 1 describes a fuzzy extractor scheme and its terminology. We will use the same structure and terminology for our proposed obfuscation method. The Gen function generates a secret key and auxiliary data $P$ from a biometric template $t$. The Rep function takes a biometric query $q$ and the auxiliary data as its input. If dist($q,t$) is smaller than a preset threshold, then it generates the correct key. In other words, the auxiliary data $P$ helps to remove noise from query $q$ in order to generate $t$. The auxiliary data $P$ is required to be secure and reveal limited information about template $t$ to preserve the privacy of template $t$. To satisfy the information-theoretic security, the entropy loss must be at most $\mathcal{L}$, as further detailed below [14], [13], [6].
@@
The rest of this paper is organized as follows: Section II reviews related work. Section III describes our methodology. Section IV describes the security capability of our scheme. Section V outlines the experimental setup. Section VI presents our results and their implications. Finally, we conclude with a summary of our method’s unique attributes, challenges faced, and future work.
@@ II Related Work
Many methods have been introduced over the years to preserve the privacy of biometric data. For instance, [32] used a de-identification technique, but anonymity was not guaranteed. Other methods explored encryption-based schemes [30], but these are vulnerable to various attacks, including those that exploit high False Acceptance Rates (FAR), where an attacker attempts to gain unauthorized access by presenting multiple fraudulent biometric samples until one is falsely accepted as genuine. Our approach, demonstrably secure against such attacks (Section IV), offers significant improvements. Techniques utilizing Convolutional Neural Networks (CNNs) have been used to protect face templates while maintaining good matching performance [21]. Blockchain-based storage of biometric templates has also been discussed [11], but it often incurs high execution times. Adversarial learning for concealing information within a learned space has also been explored [29], focusing on disentangling identity from attribute information to mitigate soft-biometric information in face templates [42, 3]. Protecting source data by combining biometric modalities (e.g., fingerprint and iris scans) using various fusion techniques has been studied [39, 35]. While enhancing security through redundancy, it introduces a single point of failure, making the entire system vulnerable if one trait is compromised. It is also possible to show the unlinkability of the protected templates using different existing measures such as maximal leakage [37]. In recent years, Generative AI (GenAI) has gained traction in biometric privacy. GenAI’s ability to generate highly realistic synthetic biometric data shows promise for enhancing privacy and security. For instance, [16] and [26] demonstrated the effectiveness of Generative Adversarial Networks (GANs) in creating synthetic facial images that can obfuscate original data, and thus protect user identities. Similarly, [48] explored using GenAI to generate diverse biometric data, improving the robustness of biometric systems against various attacks. Recent studies have investigated GenAI’s role in creating synthetic multimodal biometric data, for instance combining facial images with other modalities such as voice and fingerprints [45, 7]. This approach not only enhances security but also provides additional layers of privacy protection by making it more difficult to link synthetic data to real individuals. Furthermore, [51] highlights the potential of GenAI in generating synthetic data for training biometric systems, reducing reliance on real biometric data, and thus mitigating some privacy concerns.
One can leverage existing measures, such as maximal leakage [37], to demonstrate the unlikability of protected templates. Our approach leverages such GenAI advances to provide a robust and secure method for protecting biometric templates, addressing a number of the limitations of previous methods while being computationally efficient.
@@ III-A Main Idea
In this section, we will first informally explain the main idea behind the proposed method, and then provide the formalized algorithm. The idea is to divide the biometric template into $m$ sub-templates and hide them with $n$ chaff points to make it computationally impractical for an adversary to find the template. One solution would be to hide a template $t$ with $n$ chaff points in a public vault, $V=[r_{1},...,t,...,r_{n}]$, and store $f(t)$, where $f$ is a one-way function such as SHA-256. Now the question is how hard it would be for an adversary to find the index of $t$ in $V$. To make it computationally impractical for the adversary to find $t$, $n$ must be a very big number $(\lambda$ (security parameter)), as the best strategy to find the template $t$ is to compute $f(x)$ for all $x$ in the vault to find the template $t$. However, this method is also computationally impractical for the legitimate user. By way of example, assume we will have $f(t)$ and one vault containing $1+2^{80}$ vectors as auxiliary data. We assume the inserted chaff points are indistinguishable from the template $t$ in the collection. By indistinguishable chaff points, we mean that the chaff points represent realistic face templates that an attacker cannot differentiate from the actual template. We achieve this using StyleGAN2 to generate synthetic face embeddings that closely mimic the statistical properties of genuine templates, making it extremely challenging to identify the original among them. For matching a query $q$, we compute the distance of $q$ from all points inside the vault and retrieve $m$ (depending on the threshold of the system) closest points to $q$: ${p_{1},p_{2},...,p_{m}}$, and compute $f(p_{1}),f(p_{2}),...,f(p_{m})$ and check if there are any matches between these values and $f(t)$. This system is secure under certain conditions, but impractical because during the authentication the user needs to compute the desired distance of the query $q$ vs. $1+2^{80}$ vectors.
@@
Thus we propose generating $n$ sub-templates $t_{1},t_{2},...,t_{n}$ from template $t$, using protocol $\mathcal{G}$, such that $\mathcal{G}(t):t_{1},t_{2},...,t_{n}$ and hide each of them with $m$ chaff points. In the end, we store $f([t_{1},t_{2},...,t_{n}])$ (, where $f$ is a one-way function), in $n$ vaults such that each vault has $m+1$ vectors. To make it computationally impractical for the adversary to find the template by brute force, we need $m^{n}$ to be very large. Unlike the earlier impractical example, we can choose $m$ and $n$ to garner computational security against brute force attacks while having a practical solution. In the forthcoming experimental demonstration, we use $m=2000$, and $n=5$, so that $m^{n}=2^{54.82}$ to get acceptable security against a brute force attack. When matching query $q$, we generate $q_{1},q_{2},...q_{n}$ using protocol $\mathcal{G}(q):q_{1},q_{2},...,q_{n}$ (protocol $\mathcal{G}$ will be explained later). The distances between $q_{i}$ and all $m+1$ vectors in the corresponding vault are then computed and the top $s$ vectors from each vault are retrieved. The threshold of the system is $s$ and the distance can be computed by any distance metric, such as cosine similarity. At this point, we have retrieved $sn$ vectors. Next, we construct all $s^{n}$ possible combinations of vectors $v_{1},v_{2},...v_{s^{n}}$ and compute $f(v_{1}),f(v_{2}),...f(v_{s^{n}})$ and check if there are any matches between the stored value of $f([t_{1},t_{2},...,t_{n}])$ and the $f(v_{i})$. If $q$ is a genuine vector, it should be able to retrieve $t_{1},t_{2},...,t_{n}$ with the threshold $s$. If $q$ is an imposter, it will fail to retrieve $t_{1},t_{2},...t_{n}$ from the vaults. It should be noted that $s$ cannot be a large number as it would render the biometric matching computationally impractical.
III-B True Positive Rate (TPR) Improvement
@@ III-C The Algorithm
Sub-template Generator $\mathcal{G}(t,n)$: This protocol takes template $t$ and the number of vaults $n$ as its inputs and generates $n$ independent sub-templates $T=(t_{1},t_{2},...,t_{n})$. In the experimental section, we propose one solution for such a function.
@@ IV Security
Revocability: A biometric system needs to be revocable because if it is compromised, or if the user wishes to no longer be identified, they should be able to revoke their reference. Our scheme’s templates are revocable by choosing a new set of chaff points and a new sub-template generator. This process is further facilitated by the ease of GAN-based chaff generation.
@@
Diversity: A secure template must not allow cross-matching across databases, thereby ensuring the user’s privacy. In our system, for a false positive to occur, the imposter must meet the authentication criteria of each vault and successfully retrieve all segments of the template, hence our system mitigates the cross-matching as compared to systems using one model. It also could be verified throughout the experimental section, the FPR is very small, because an imposter needs to satisfy multiple vaults to be able to positively match.
@@
Accuracy: Our method shows minimum accuracy degradation with a 0.99 area under the curve (AUC) for both the AT&T  [4] and Georgia Tech (GT) face databases [31].
@@
Security: To show that our system is secure, we need to show that the auxiliary data $P=(P_{1},P_{2})$ is not revealing too much information and there is enough entropy in this system. Chaff points are generated from the same distribution that the corresponding sub-template is generated. This is achieved using StyleGAN2, which was originally trained on the Flickr-Faces-HQ (FFHQ) [24]dataset containing 70,000 high-quality, diverse human face images. This training ensures that the generated chaff points closely mimic the statistical properties and distribution of real face embeddings. Hence we can say the sub-templates inside each vault are indistinguishable from other chaff points there. Getting this indistinguishability between the sub-template and chaff points inside a vault is straightforward; the challenge is to achieve indistinguishability between sub-templates and chaff points. In other words, if an adversary knows that $t_{1}$ is a sub-template, other $t_{i}$ should remain indistinguishable from chaff points, we achieve this by using independent models, given two models $M_{1},M_{2}$, and template $t$ and a random face $r$, let ======= = = be coming from the same image.
$|Pr[M_{1}(t)\text{===}M_{2}(t)]-Pr[M_{1}(t)\text{===}M_{2}(r)]|=\epsilon$
@@
To achieve this, we must make sure that the sub-templates are independent. If we assume everything looks completely random inside vaults, then we can claim the best strategy to find the template for any adversary is to check all possible combinations of elements in vaults, which we have made impractical by choosing proper $n$,$m$. As a result, our approach provides superior security against brute force attacks, as well as faster face acceptance and rejection capabilities.
@@
It is clear that given a face template $t$, a random face $r$, and two biometric models $M1$ and $M2$, $y_{1}=M1(t)$ and $y_{2}~{}=~{}M2(t)$ cannot be completely independent. However, in our case, we are using big neural networks. With this assumption and if $M1$ and $M2$ have completely different structures, it is not going to be easy to distinguish between $M1(t)$ and $M1(r)$ given $M2(t)$. In other words, if we tell the adversary $M2(t)$ belongs to person $A$, and then if we ask him to tell us which one of $M1(t)$ and $M1(r)$ belongs to the same person, it will not be easy for the adversary to solve this problem.
@@
One approach is to reverse engineer the template to recreate it and see which one of reverse $M1(t)$ and reverse $M1(r)$ is closer to the reverse $M2(t)$. We have addressed this attack by adding randomness; now the adversary needs to distinguish between $R_{2}M1(t)+R_{2}^{\prime}$ and $R_{2}M1(r)+R_{2}^{\prime\prime}$, given $R_{1}M2(t)+R_{1}^{\prime}$.
V Experiments
V-A Experimental Setup